TGBT18B_502_02 - Incoherent BOF files check outs


  • There was a bug in a previous version of incoherent dedispersion BOF files causing an 8-channel shift in the spectrum. Goal here is to test new versions with a test-tone injection to confirm this has been fixed, before moving on to astronomical check-out.


  • Session begins at 15:35. Need to find someone to stick me in the gateway. First scan at 15:51
  • Note: Source name is left over and is "JUPITER". Just ignore this.
  • Testing i1500x1024 first with i1500x1024_x14_7_24t_095binw_2018_Oct_05_1350.bof
    • Didn't see test tone in first scan. Sampling rate was left over from a coherent set up so it's possible there as lots of dropped data. Running again (also turning scales down to 10000...were at 20000)
    • Looks better on second scan, but there is RFI at the frequency I chose (1670 MHz). Increasing tone power and moving to 1750 MHz for tone frequency.
    • Tone is saturating but is there. It seems like there might be a small offset, but hard to tell if it is just a resolution issue. Will move on to next BOF and see what that looks like.
  • Testing i1500x2048 with i1500x2048_x14_7_24t_095binw_2018_Sep_27_1307.bof
    • Tone looks good. Again, small offset but it might just be a resolution issue. Shifting the frequency so that it falls in the middle of a channel.
    • Tone seems to be split across a couple of channels. Might need to think of a way to get a more precise measurement. But will move on for now.
  • Testing i0800x1024 with i0800x1024_x14_7_24t_095binw_2018_Oct_15_1559.bof
    • Tone is back at 1750 MHz. Seems to be right on the money.
  • Testing i0800x2048 with i0800x2048_x14_7_24t_095binw_2018_Feb_15_1118.bof
    • Also looks good.
  • Testing i0800x0064 with i0800x0064_x14_7_24t_095binw_2018_Oct_16_1222.bof
    • Looks good.
  • Testing i0800x0128 with i0800x0064_x14_7_24t_095binw_2018_Oct_16_1222.bof
    • Looks good.
  • Testing i0800x0256 with i0800x0256_x14_7_24t_095binw_2018_Oct_16_0654.bof
    • Looks good
  • Testing i1500x0064 with i1500x0064_x14_7_24t_095binw_2018_Oct_12_1300.bof
    • Changed tone to 1757.8125. Looks right on.
  • Testing i1500x0128 with i1500x0128_x14_7_24t_095binw_2018_Oct_12_0759.bof
    • Also looks good.
  • Testing i1500x0256 with i1500x0256_x14_7_24t_095binw_2018_Oct_10_1521.bof
    • Looks good
  • Scans 14 and 15 are i1500x1024 and i1500x2048 (same BOFs as above) but with tone injection at 1757.8125.
    • Both look right on now. Apparent shift was just a resolution and plotting issue.
  • Testing i1500x4096 with i1500x4096_x14_7_24t_095binw_2018_Sep_26_0724.bof. Using acclen of 64
    • Scan is #17 (stopped 16 since I forgot to save the SB). Right on.
  • i0800x4096 with i0800x4096_x14_7_24t_095binw_2018_Sep_19_0906.bof
    • Tone is at 1750 MHz. Data seem to have been written but astropy is complaining that there is no data extension in the last HDU. psrplot is also unhappy about the number of subintegrations. Running a second scan.
    • No issues on second time. Tone right where it should be.
  • Trying C-Band setup from 17A-433 to check variances. Search mode, i1500x1024, four banks. No test tone.
    • VEGAS never started taking data on first scan. Not sure why. Had an abort on next attempt. Did a VEGAS Off/On cycle and it seems to be OK now.
  • Switched BOF files back. VEGAS had some issues coming back up. Not sure why. Seemed like it was taking a long time to activate and program the ROACH's. Seems OK now.
  • Did a Mode 1 checkout as well. Looks good.


  • All these BOF files seems OK. Will do astronomical commissioning when possible.

-- RyanLynch - 2018-12-20
Topic revision: r1 - 2018-12-20, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback